OpenStudy (anonymous):

Gwen has 3 red erasers, 4 green erasers, 2 black erasers, 5 purple erasers, and 6 yellow erasers in her book bag. What is the probability Gwen will reach in her book bag and select a purple eraser? is it 1/4?

OpenStudy (anonymous):

looks like it

OpenStudy (anonymous):

the total erasers are 3+4+2+5+6 = 20 p( purple) = 5/20 = 1/4 you are correct

OpenStudy (anonymous):

thanks